OpenStudy (anonymous):

For the function f(x) = (8-2x)2 ,find f-1 . Determine whether f-1 is a function.

OpenStudy (agent0smith):

\[\Large y = (8-2x)^2\] solve for x start by taking square roots of both sides

OpenStudy (anonymous):

sqrts of both sides?

OpenStudy (jdoe0001):

\(\bf \sqrt{y}=\sqrt{(8-2x)^2}\leftarrow that\)

OpenStudy (anonymous):

And what do i do with that?

OpenStudy (agent0smith):

Square root cancels out a square

OpenStudy (anonymous):

i dont get it ?

OpenStudy (agent0smith):

\[\large \sqrt {x^2} = x\] \[\large \sqrt{(x-6)^2} = x-6\] etc

OpenStudy (anonymous):

oh okay so it would be the same thing is was before it was sqrt with out the 2?

OpenStudy (agent0smith):

Er, not sure what you mean, just write it.

OpenStudy (anonymous):

y = (8-2x)

OpenStudy (agent0smith):

Sqrt y is still sqrt y.

OpenStudy (anonymous):

oh so its sqrt y= (8-2x) ?

OpenStudy (jdoe0001):

\(\bf \sqrt{y}=\sqrt{(8-2x)^2} \\ \quad \\ \sqrt[{\color{red}{ 2}}]{x^{\color{red}{ 2}}} \iff x\qquad thus \\ \quad \\ \sqrt{y}=\sqrt[{\color{red}{ 2}}]{(8-2x)^{\color{red}{ 2}}}\implies \sqrt{y}=8-2x\)

OpenStudy (agent0smith):

Now subtract 8 from both sides then divide by -2

OpenStudy (anonymous):

sqrty -8= -2x?

OpenStudy (agent0smith):

Yes

OpenStudy (anonymous):

now what?

OpenStudy (agent0smith):

See my last post, it's there

OpenStudy (anonymous):

Which one?

OpenStudy (agent0smith):

Now subtract 8 from both sides then divide by -2

OpenStudy (anonymous):

x=sqrt y-8/ -2 ?

OpenStudy (agent0smith):

you forgot parentheses :P \[\large x =\frac{ ( \sqrt y -8 ) }{-2 }\]

OpenStudy (agent0smith):

Simplify, by multiplying the top and bottom of the fraction by -1: \[\large x =\frac{ 8- \sqrt y }{2 }\]now just swap x and y.

OpenStudy (anonymous):

y= 8- sqrt x/2 ?

OpenStudy (anonymous):

Its not a function right?

OpenStudy (agent0smith):

It's not because we forgot to put the plus or minus in, which you have to do when taking square roots and you forgot parentheses again :P

OpenStudy (agent0smith):

\[\large x =\frac{ 8\pm \sqrt y }{2 }\]

OpenStudy (anonymous):

Okay :) thanks :)